# Equations involving the modular $j$-function and its derivatives

### v1

@@ 5.1. Generic transforms
By construction, for any $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, if $n=\deg_{X}(p)$ we have
@@
The map $\Gamma$ is a $K$-algebra homomorphism with left inverse $p=\Gamma(p)(X,1,0,\mathbf{Y})$ and $p\mapsto p^{\gamma}$ is multiplicative. Moreover, for any two $\gamma,\gamma^{\prime}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, we have $(p^{\gamma})^{\gamma^{\prime}}=r(X)p^{\gamma^{\prime}\gamma}$ for some $r\in\operatorname{\mathbb{Z}}[X]$, hence $p$ is irreducible over $\operatorname{\mathbb{C}}(X)$ if and only if so is $p^{\gamma}$.
@@
Suppose that $p(z,\mathbf{j}(z))=0$ has a Zariski dense set of solutions and suppose that all the solutions of $h(z,\mathbf{j}(z))=0$ also satisfy $r(z,\mathbf{j}(z))=0$. Then the solutions of the original equation satisfy $r^{\gamma^{-1}}(z,\mathbf{j}(z))=0$, thus in particular $p$ divides $r^{\gamma^{-1}}$, and so $h$ divides $(r^{\gamma^{-1}})^{\gamma}\in r\operatorname{\mathbb{Z}}[X]$, thus it divides $r$, showing that the solutions of $h(z,\mathbf{j}(z))=0$ are Zariski dense. ∎
@@ 5.2. Density criteria
Then there are $\ell>0$, $10\leq x_{0}<$ 1, $M>0$, $E_{0}>0$ such that for all $m\in\operatorname{\mathbb{Z}}$ sufficiently large the function $F(z,\mathbf{j}(z))$ has $\ell$ zeroes, counted with multiplicity, within the region $m+\gamma\Xi_{m}$, where
@@
Likewise, set $G(z)\coloneqq F(z,\mathbf{j}(z))$. By Proposition 4.6 applied to $F(z,\mathbf{j}(\gamma w))$, $F_{m}(z)$ has $\ell>0$ many zeroes in a certain region $\Xi_{m}$ and is suitably bounded from below for $z\in\partial\Xi_{m}$, as long as $m$ is sufficiently large.
If $\gamma$ is upper triangular, namely $\gamma z=z+k$ for some $k$, we choose $s=1$, $t=0$, and observe that trivially
@@
Thus, as soon as $z$ is sufficiently large, the last factor on the right hand side has modulus less than $\frac{1}{2}$ independently of $m$. Then pick $m$ large enough so that this happens whenever $\operatorname{Im}(z)>E_{0}\log|m|$, and so
@@
Therefore, by Rouché’s theorem, $G(m+\gamma z)$ and $F_{m}(z)$ have the same number of zeroes in $\Xi_{m}$, counted with multiplicity. It follows that $G(z)$ has $\ell$ zeroes in the region $m+\gamma\Xi_{m}$. ∎
@@
Suppose by contradiction that all the solutions of $F(z,\mathbf{j}(z))=0$ lie on a further hypersurface $G=0$, where $G$ is a non-constant polynomial not divisible by $F$. In particular, the algebraic set defined by $\{F=G=0\}$ has dimension two, so its projection onto the variables $Y_{0},Y_{1},Y_{2}$ has dimension at most two, meaning that the solutions satisfy an equation $H(\mathbf{j}(z))=0$ for some non-constant polynomial $H$. The assumption on $F$ implies that $F$ depends on the variable $X$, thus $F$ and $H$ are coprime.
By Propositions 4.1 and  5.3, for $m\in\operatorname{\mathbb{Z}}$ large, there are regions $E_{m}$ such that the original equation has solutions in $m+E_{m}$, and moreover the real part of each $E_{m}$ is bounded from above and below and the imaginary part is bounded from below uniformly in $m$. Each solution can be in $m+E_{m}$ for at most finitely many $m\in\operatorname{\mathbb{Z}}$. This implies that for some $m$ sufficiently large, the function $H(\mathbf{j}(z))$ has infinitely many zeroes in the region $\bigcup_{|k|>|m|}E_{k}$, and so it is constantly zero (trivially if that union is bounded, and by Lemma 4.3 otherwise), contradicting the algebraic independence of $j$, $j^{\prime}$, $j^{\prime\prime}$. ∎
@@
Let $F(X,\mathbf{Y})=\sum_{k=0}^{n}X^{k}p_{k}(\mathbf{Y})$ be irreducible. Suppose that $p_{n}$ has a factor $h$ such that $h(\mathbf{j}(z))=0$ has a Zariski dense set of solutions. Then the equation $F(z,\mathbf{j}(z))=0$ has a Zariski dense set of solutions.
@@ 6. Zero Estimates
Given $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$, we will prove a few estimates on the order of $p(z,\mathbf{j}(z))$ at different points, distinguishing three cases. Before that, we note that specialising the variables $Y_{1}$ and $Y_{2}$ of $\Gamma(p)$ will almost always return a simple twist of the original polynomial, in the following sense.
@@
Let $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ be non-zero, $\tau\in\operatorname{\mathbb{H}}$. Suppose that $j^{\prime}(\tau)\neq 0$ and let $s$ be the maximum integer such that $(Y_{0}-j(\tau))^{s}$ divides $p$. Then for all $\gamma$ in a Zariski dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the function $p(z,\mathbf{j}(z))$ has order $s$ at $z=\gamma\tau$.
@@
Clearly, it suffices to prove the conclusion when $s=0$, namely when $p$ is not divisible by $Y_{0}-j(\tau)$. By Lemma 6.1 at $\alpha^{2}=j^{\prime}(\tau)\neq 0$, $\beta=j^{\prime\prime}(\tau)$, $U_{1}^{2}=Y_{1}$, $U_{2}=Y_{2}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
Since $p$ is not divisible by $Y_{0}-j(\tau)$, $p(X,j(\tau),Y_{1},Y_{2})$ is non-trivial, hence in particular $r(Z,W,C)\coloneqq\Gamma(p)(Z,W,C,\mathbf{j}(\tau))$ is non-trivial.
@@
$p^{\gamma}(\tau,\mathbf{j}(\tau))=(c\tau+d)^{n}r(\gamma\tau,c\tau+d,c)=(c\tau+d)^{n}p(\gamma\tau,\mathbf{j}(\gamma\tau)).$
In particular, $p(\gamma\tau,\mathbf{j}(\gamma\tau))=0$ if and only if $r(\gamma\tau,c\tau+d,c)=0$. Since the map $\gamma\mapsto(\gamma\tau,c\tau+d,c)$ from $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$ to $\operatorname{\mathbb{C}}^{3}$ is injective, hence dominant, and $r$ is non-trivial, this implies that for $\gamma$ outside of some proper Zariski closed set, $p(z,\mathbf{j}(z))$ has order $s=0$ at $z=\gamma\tau$. ∎
With this proposition, we immediately obtain the following special case of Theorem 1.2.
@@
The equation $j(z)-u=0$ has a Zariski dense set of solutions if and only if $u\notin\{0,1728\}$.
Now we consider the behaviour of $p(z,\mathbf{j}(z))$ towards the cusps of the fundamental domains. One can easily see that the order at the cusp of any given fundamental domain is at least $(-e,-N)$ for some $N$, where $e=\deg_{T}(p(X,T\mathbf{Y}))$. This is not far from the actual behaviour at most cusps.
@@
Let $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$. Then there is $0\leq M\leq\deg_{W}(\Gamma(p))$ such that for all $\gamma$ in a Zariski dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the function $p(z,\mathbf{j}(z))$ has order $(-e,-M)$ at the cusp of $\gamma\mathbb{F}$, where $e$ is the degree of $p(X,T\mathbf{Y})$ in $T$.
@@
Since obviously $\Gamma(p)(Z,W,C,T\mathbf{Y})=\Gamma(p(X,T\mathbf{Y}))$, we have
@@
Therefore, the order of $p^{\gamma}(z,\mathbf{j}(z))$ at $i\infty$ is $(-e,-(n+M))$, and so the order of $p(\gamma z,\mathbf{j}(\gamma z))$ is $(-e,-M)$, unless the leading coefficient of $r(\nicefrac{{a}}{{c}},W,c)$ vanishes or $c=0$.
To conclude, since the map $\gamma\mapsto(a,c)$ from $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$ to $\operatorname{\mathbb{C}}^{2}$ is dominant, it suffices to prove that $r$ is non-trivial. By Lemma 6.1 at $\alpha^{2}=-2\pi i$, $\beta=-4\pi$, $U_{1}^{2}=Y_{1}Y_{0}^{-1}$, $U_{2}=Y_{2}Y_{0}^{-1}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
which has order $(-1,-4)$ at $i\infty$, since $z^{4}j^{\prime\prime}(z)$ is obviously the dominant term. Note that here $4=\deg_{W}(\Gamma(p))$.
@@
has order $(-2,-3)$ for $c\neq 0$, because $j(z)j^{\prime\prime}(z)\sim(j^{\prime}(z))^{2}\sim-4\pi^{2}q^{-2}$, and it obviously has order $(-2,0)$ for $c=0$. On the other hand, $\deg_{W}(\Gamma(h))=4$. Proposition 6.4 guarantees that even though these cancellations may occur at all cusps, some term of maximal exponential growth is not cancelled, at least generically.
@@
Let $p\in\operatorname{\mathbb{C}}[X,Y_{0},Y_{1}]$, $\tau\in\operatorname{\mathbb{H}}$, $u=j(\tau)$, and $\mu$ be the order of $j(z)-u$ at $z=\tau$. Suppose that $\mu>1$. Then for all $\gamma$ in a Zariski dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the order of $p(z,\mathbf{j}(z))$ at $z=\gamma\tau$ is the highest power of $T$ dividing $p(X,T^{\mu}Y_{0}+u,T^{\mu-1}Y_{1})$.
@@
The proof is very similar to that of Proposition 6.4. Write
@@
It follows that the order of $p^{\gamma}(z,\mathbf{j}(z))$ at $z=\tau$ is $\nu$ as long as $r$ and $c$ do not vanish. Since $z\mapsto\gamma z$ is a diffeomorphism, this coincides with the order of $p(z,\mathbf{j}(z))$ at $z=\gamma\tau$.
To conclude, since the map $\gamma\mapsto(\gamma\tau,c\tau+d,c)$ is injective on $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$, hence dominant on $\operatorname{\mathbb{C}}^{3}$, we only need to show that $r$ is non-trivial. Pick $\beta_{0}$ such that $\beta_{0}^{2\mu}=\alpha_{0}^{-1}$ and $U_{0}$ such that $U_{0}^{2\mu}=Y_{0}$. By Lemma 6.1 at $\alpha^{2}=\mu\alpha_{0}$, $\beta=0$, $U_{1}^{2}=(\beta_{0}U_{0})^{-2(\mu-1)}Y_{1}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
$p(\mathbf{j}(z))=\alpha^{2}(z-\gamma\rho)^{4}\left(6-\frac{2}{3}3^{2}+o(1)\right)=o((z-\gamma\rho)^{4}).$
@@
is non-constant, then it has a pole in $\operatorname{\mathbb{H}}$ or exponential growth on some fundamental domains.
@@
Suppose that $P(z)$ has no poles in $\operatorname{\mathbb{H}}$ and has no exponential growth in any fundamental domain. Without loss of generality, we may assume that $p$ and $h$ are coprime.
Let $\beta$ be a root of $h(Y_{0})$ such that $\beta\notin\{0,1728\}$, and let $\tau$ be such that $j(\tau)=\beta$. In particular, $Y_{0}-\beta$ does not divide $p$. By Proposition 6.2, $p(\gamma\tau,j(\gamma\tau))\neq 0$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ except for some proper Zariski closed set, and so $P$ has a pole at $\gamma\tau$, a contradiction. Therefore, $h(Y_{0})$ is of the form $\alpha Y_{0}^{s}(Y_{0}-1728)^{t}$ for some $\alpha\in\operatorname{\mathbb{C}}$.
Now $h(j(z))j^{\prime}(z)^{\ell}$ has order $3s+2\ell$ at all the conjugates of $\rho$ and $2t+\ell$ at all the conjugates of $i$, thus $p(j(z),j^{\prime}(z))$ must have at least the same order at those points. On writing $p=\sum_{\ell^{\prime}}Y_{1}^{\ell^{\prime}}p_{\ell^{\prime}}(Y_{0})$, Proposition 6.6 implies that each $p_{\ell^{\prime}}(Y_{0})$ is divisible by $Y_{0}^{s+\lceil\frac{2(\ell-\ell^{\prime})}{3}\rceil}$ and by $(Y_{0}-1728)^{t+\lceil\frac{\ell-\ell^{\prime}}{2}\rceil}$. In particular, whenever $p_{\ell^{\prime}}\neq 0$ we have

### v3

@@ 5.1. Generic transforms
In particular $\deg_{X}(p)=\deg_{Z}(\Gamma(p))$. By construction, for any $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, if $n=\deg_{X}(p)$ we have
@@
We make the following observations.
  1. O1.
    The map $\Gamma:K[X,\mathbf{Y}]\to K[Z,W,C,\mathbf{Y}]$ defined above is a $K$-algebra homomorphism with left inverse $p(X,\mathbf{Y})=\Gamma(p)(X,1,0,\mathbf{Y})$.
  2. O2.
    For each $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the map $p\mapsto p^{\gamma}$ is multiplicative, that is $(p_{1}p_{2})^{\gamma}=p_{1}^{\gamma}p_{2}^{\gamma}$ for any $p_{1},p_{2}\in K[X,\mathbf{Y}]$. Indeed, this follows at once from the fact that $\Gamma$ is a homomorphism (O1) and $\deg_{X}(p_{1}p_{2})=\deg_{X}(p_{1})+\deg_{X}(p_{2})$.
  3. O3.
    For any $p\in K[X,\mathbf{Y}]$ and any $\gamma_{1},\gamma_{2}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ there is $r\in\operatorname{\mathbb{Z}}[X]$ such that $(p^{\gamma_{1}})^{\gamma_{2}}=r(X)p^{\gamma_{1}\gamma_{2}}$.
    Indeed, let $n=\deg_{X}(p)$ and $m=\deg_{X}(p^{\gamma_{1}})$. By construction, we have $m=n+\deg_{Y_{0}}(p)\geq n$. Now write $\gamma_{t}=\begin{pmatrix}a_{t}&b_{t}\\ c_{t}&d_{t}\end{pmatrix}$, for $t\in\{1,2\}$, and write $\gamma_{1}\gamma_{2}=\begin{pmatrix}\widetilde{a}&\widetilde{b}\\ \widetilde{c}&\widetilde{d}\end{pmatrix}$. Thus
    $\frac{p^{\gamma_{1}\gamma_{2}}(z,\mathbf{j}(z))}{\left(\widetilde{c}z+\widetilde{d}\right)^{n}}=p(\gamma_{1}\gamma_{2}z,\mathbf{j}(\gamma_{1}\gamma_{2}z))=\frac{p^{\gamma_{1}}(\gamma_{2}z,\mathbf{j}(\gamma_{2}z))}{(c_{1}\gamma_{2}z+d_{1})^{n}}=\frac{\left(p^{\gamma_{1}}\right)^{\gamma_{2}}(z,\mathbf{j}(z))}{(c_{1}\gamma_{2}z+d_{1})^{n}(c_{2}z+d_{2})^{m}}.$
    Since $m\geq n$ and $(c_{1}\gamma_{2}z+d_{1})(c_{2}z+d_{2})=\widetilde{c}z+\tilde{d}$, we get
    $r(X):=\frac{(c_{2}X+d_{2})^{m}}{\left(\widetilde{c}X+\widetilde{d}\right)^{n}}\left(c_{1}\frac{a_{2}X+b_{2}}{c_{2}X+d_{2}}+d_{2}\right)^{n}=(c_{2}X+d_{2})^{m-n}\in\operatorname{\mathbb{Z}}[X].$
  4. O4.
    For any $p\in K[X,\mathbf{Y}]$ and any $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, if we consider $p$ and $p^{\gamma}$ as polynomials in the variables $\mathbf{Y}$ with coefficients in $K(X)$, then $p$ is irreducible in $K(X)[\mathbf{Y}]$ if and only if so is $p^{\gamma}$.
    Indeed, note that if $p$ is not a unit (meaning it contains one of the variables $Y_{0}$, $Y_{1}$, $Y_{2}$), then $p^{\gamma}$ is also not a unit. It follows by (O2) that if $p$ is reducible in $K(X)[\mathbf{Y}]$, thus it is a product of two non-units, then so is $p^{\gamma}$. Likewise, if $p^{\gamma}$ is reducible, then $(p^{\gamma})^{\gamma^{-1}}$ is reducible too, and $(p^{\gamma})^{\gamma^{-1}}=r(X)p$ for some $r(X)\in\operatorname{\mathbb{C}}[X]$ by (O3); since $r(X)$ is a unit, it follows that $p$ is reducible.
@@
By (O3) and (O4), $p^{\gamma}=r(X)h$ and $h^{\gamma^{-1}}=s(X)p$ for some $r,s\in\operatorname{\mathbb{C}}[X]$. It follows that $p(z,\mathbf{j}(z))=0$ and $h(z,\mathbf{j}(z))=0$ have the same solutions except possibly for the zeroes of $r(z)$ and $s(z)$. Since those are only finitely many, the solutions of the former equation are Zariski dense if and only if so are the solutions of the latter. ∎
@@ 5.2. Density criteria
Then there are $\ell>0$, $0\leq x_{0}<1$, $M>0$, $E_{0}>0$ such that for all $m\in\operatorname{\mathbb{Z}}$ sufficiently large the function $F(z,\mathbf{j}(z))$ has $\ell$ zeroes, counted with multiplicity, within the region $m+\gamma\Xi_{m}$, where
@@
Likewise, set $G(z)\coloneqq F(z,\mathbf{j}(z))$. By Proposition˜4.8 applied to $F(z,\mathbf{j}(\gamma z))$, $F_{m}(z)$ has $\ell>0$ zeroes in a certain region $\Xi_{m}$ and is suitably bounded from below for $z\in\partial\Xi_{m}$, as long as $m$ is sufficiently large.
If $\gamma$ is upper triangular, namely $\gamma z=z+k$ for some $k$, we choose $s=1$, $t=0$, and observe that since the functions of $\mathbf{j}$ are 1-periodic
@@
Thus, as soon as $z$ is sufficiently large, the last factor on the right-hand side has modulus less than $\frac{1}{2}$ independently of $m$. Then pick $m$ large enough so that this happens whenever $\operatorname{Im}(z)>E_{0}\log|m|$, and so
@@
Therefore, by Rouché’s theorem 3.1, $G(m+\gamma z)$ and $F_{m}(z)$ have the same number of zeroes in $\Xi_{m}$, counted with multiplicity. It follows that $G(z)$ has $\ell$ zeroes in the region $m+\gamma\Xi_{m}$. ∎
@@
Suppose by contradiction that all the solutions of $F(z,\mathbf{j}(z))=0$ lie on a further hypersurface $G=0$, where $G\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ is a non-constant polynomial not divisible by $F$. In particular, the algebraic subset of $\operatorname{\mathbb{C}}^{4}$ defined by $\{F=G=0\}$ has dimension two, so its projection onto the variables $Y_{0},Y_{1},Y_{2}$ has dimension at most two, meaning that the solutions satisfy an equation $H(\mathbf{j}(z))=0$ for some non-constant polynomial $H\in\operatorname{\mathbb{C}}[\mathbf{Y}]$. The assumption on $F$ implies that $F$ depends on the variable $X$ (i.e. $n\geq 1$), thus $F$ and $H$ are coprime.
By Propositions 4.1 and  5.3, for $m\in\operatorname{\mathbb{Z}}$ large, there are regions $\Xi_{m}$ such that the original equation has solutions in $m+\Xi_{m}$, and moreover the real part of each $\Xi_{m}$ is bounded from above and below and the imaginary part is bounded away from 0 uniformly in $m$. For each solution $\tau\in m+\Xi_{m}$, there are only finitely many integers $k$ such that $\tau+k$ is also a solution. This implies that for some $m$ sufficiently large, the function $H(\mathbf{j}(z))$ has infinitely many zeroes in the region $\bigcup_{|k|>|m|}\Xi_{k}$. If this union is bounded, then we conclude that $H$ is constantly zero by the identity theorem from complex analysis, but this contradicts the algebraic independence of $j$, $^{\prime}, j′′j^{\prime\prime}$. So we assume that the union is unbounded, but by Lemma˜4.3 there exist $\alpha\in\operatorname{\mathbb{C}}^{\times}$, $d,e\in\operatorname{\mathbb{Z}}$ and $C>0$ such that $H(z)\sim\alpha z^{d}\exp(e2\pi iz)$ in the region $U:=\{z\in\operatorname{\mathbb{H}}:\mathrm{Im}(z)\geq C\log|z|\}$. Then the only way $H$ can have infinitely many zeroes in $\bigcup_{|k|>|m|}\Xi_{k}$ is if $H$ is constantly zero, again contradicting the algebraic independence of $j$, $^{\prime}, j′′j^{\prime\prime}$. This completes the proof. ∎
@@
Let $F(X,\mathbf{Y})=\sum_{k=0}^{n}X^{k}p_{k}(\mathbf{Y})$ be irreducible. Suppose that $p_{n}$ has a factor $h$ such that the equation $h(\mathbf{j}(z))=0$ has a Zariski dense set of solutions. Then the equation $F(z,\mathbf{j}(z))=0$ has a Zariski dense set of solutions.
@@ 6. Zero Estimates
Given $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$, we will prove a few estimates on the order of $p(z,\mathbf{j}(z))$ at different points, distinguishing three cases. Before that, we note that specialising the variables $Y_{1}$ and $Y_{2}$ of $\Gamma(p)$ at some complex values will almost always return an ‘obfuscated’ copy of the original polynomial, which for instance cannot be constant unless $p$ itself was. More precisely, we note the following trivial identity.
@@
Let $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ be non-zero, $\tau\in\operatorname{\mathbb{H}}$. Suppose that $j^{\prime}(\tau)\neq 0$ and let $s$ be the maximum integer such that $(Y_{0}-j(\tau))^{s}$ divides $p$. Then for all $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the function $p(z,\mathbf{j}(z))$ has order $s$ at $z=\gamma\tau$.
@@
Let $f\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ be such that $p=(Y_{0}-j(\tau))^{s}f$; we need to show that $f(z,\mathbf{j}(z))$ has order 0 at $z=\gamma\tau$, for all $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. In other words, it suffices to prove the proposition for the case when $p$ is not divisible by $Y_{0}-j(\tau)$ (and hence $s=0$).
From now on we assume $s=0$. By Lemma˜6.1 at $\alpha^{2}=j^{\prime}(\tau)\neq 0$, $\beta=j^{\prime\prime}(\tau)$, $U_{1}^{2}=Y_{1}$, $U_{2}=Y_{2}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
Since $p$ is not divisible by $Y_{0}-j(\tau)$, $p(X,j(\tau),Y_{1},Y_{2})$ is a non-zero polynomial, hence in particular $r(Z,W,C)\coloneqq\Gamma(p)(Z,W,C,\mathbf{j}(\tau))$ is also non-zero.
@@
$p^{\gamma}(\tau,\mathbf{j}(\tau))=(c\tau+d)^{n}r(\gamma\tau,c\tau+d,c)=(c\tau+d)^{n}p(\gamma\tau,\mathbf{j}(\gamma\tau)),$
where $n=\deg_{X}(p)$. In particular, $p(\gamma\tau,\mathbf{j}(\gamma\tau))=0$ if and only if $r(\gamma\tau,c\tau+d,c)=0$.
The map $\upsilon:\operatorname{SL}_{2}(\operatorname{\mathbb{C}})\to\operatorname{\mathbb{C}}^{3}$ given by $\gamma\mapsto(\gamma\tau,c\tau+d,c)$ is injective, and since $\dim\operatorname{SL}_{2}(\operatorname{\mathbb{C}})=3=\dim\operatorname{\mathbb{C}}^{3}$, by the fibre-dimension theorem $\upsilon$ is also dominant. As $r$ is a non-zero polynomial, there is a non-empty Zariski open subset $U_{1}$ of $\operatorname{\mathbb{C}}^{3}$ such that $r$ never vanishes on $U$. This then gives a Zariski open subset $U_{0}$ of $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$ such that for every $\gamma\in U_{0}$, $\upsilon(\gamma)\in U_{1}$. This implies that for $\gamma\in U_{0}$, $p(z,\mathbf{j}(z))$ does not vanish at $z=\gamma\tau$, and hence $p(z,\mathbf{j}(z))$ has order $s=0$ at $z=\gamma\tau$. ∎
With this proposition we obtain the following special case of Theorem˜1.2.
@@
The equation j(z)u=0j(z)-u=0 has a Zariski dense set of solutions if and only if $u\notin\{0,1728\}$.
Proof.
If $u\notin\{0,1728\}$, then for any $\tau\in\operatorname{\mathbb{H}}$ satisfying $j(\tau)=u$ we have that $j^{\prime}(\tau)\neq 0$. We need to show that the solutions of $j(z)=u$ are Zariski dense. So suppose that $p(X,\mathbf{Y})\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ is such that its zero locus contains all the solutions of $j(z)=u$. The solutions of $j(z)=u$ are precisely $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\tau$, where $j(\tau)=u$, so $p(\gamma\tau,\mathbf{j}(\gamma\tau))=0$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. Therefore $p(z,\mathbf{j}(z))$ has positive order at $\gamma\tau$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, so by Proposition˜6.2, $Y_{0}-u$ divides $p$, thus proving Zariski density.
On the other hand, if $u\in\{0,1728\}$, then for any $\tau\in\operatorname{\mathbb{H}}$ satisfying $j(\tau)=u$ we have that $j^{\prime}(\tau)=0$, so the solutions of $j(z)=u$ lie in the proper Zariski closed subset given by $Y_{1}=0$. ∎
Now we consider the behaviour of $p(z,\mathbf{j}(z))$ towards the cusps of the fundamental domains. We write $T\mathbf{Y}:=(TY_{0},TY_{1},TY_{2})$. One can easily see that the order at the cusp of any given fundamental domain is at least (e,N)(-e,-N) for some $N$, where $e=\deg_{T}(p(X,T\mathbf{Y}))$. This is not far from the actual behaviour at most cusps.
@@
Let $p\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$. Then there is $0\leq M\leq\deg_{W}(\Gamma(p))$ such that for all $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the function $p(z,\mathbf{j}(z))$ has order (e,M)(-e,-M) at the cusp of $\gamma\mathbb{F}$, where $e$ is the degree of $p(X,T\mathbf{Y})$ in $T$.
@@
Since $\Gamma(p)(Z,W,C,T\mathbf{Y})=\Gamma(p(X,T\mathbf{Y}))$, we have
@@
Therefore, the order of $p^{\gamma}(z,\mathbf{j}(z))$ at $i\infty$ is (e,(n+M))(-e,-(n+M)), and so the order of $p(\gamma z,\mathbf{j}(\gamma z))$ is (e,M)(-e,-M), unless the leading coefficient of $r(\nicefrac{{a}}{{c}},W,c)$ vanishes or $c=0$.
To conclude, since the map $\gamma\mapsto(a,c)$ from $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$ to $\operatorname{\mathbb{C}}^{2}$ is dominant, it suffices to prove that $r$ is non-trivial (like we did in the proof of Proposition˜6.2). By Lemma˜6.1 at $\alpha^{2}=-2\pi i$, $\beta=-4\pi$, $U_{1}^{2}=Y_{1}Y_{0}^{-1}$, $U_{2}=Y_{2}Y_{0}^{-1}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
which has order (1,4)(-1,-4) at $i\infty$, since $z^{4}j^{\prime\prime}(z)$ is the dominant term. Note that here $4=\deg_{W}(\Gamma(p))$.
@@
has order (2,3)(-2,-3) for $c\neq 0$, because $j(z)j^{\prime\prime}(z)\sim(j^{\prime}(z))^{2}\sim-4\pi^{2}q^{-2}$, and it has order (2,0)(-2,0) for $c=0$. On the other hand, $\deg_{W}(\Gamma(h))=4$. Proposition˜6.4 guarantees that even though these cancellations may occur at all cusps, some term of maximal exponential growth is not cancelled, at least generically.
@@
Let $p\in\operatorname{\mathbb{C}}[X,Y_{0},Y_{1}]$, $\tau\in\operatorname{\mathbb{H}}$, $u=j(\tau)$, and let $\mu$ be the order of j(z)uj(z)-u at $z=\tau$. Then for all $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the order of $p(z,\mathbf{j}(z))$ at $z=\gamma\tau$ is the highest power of $T$ dividing $p(X,T^{\mu}Y_{0}+u,T^{\mu-1}Y_{1})$.
@@
The proof is very similar to that of Proposition˜6.4. Write
@@
It follows that the order of $p^{\gamma}(z,\mathbf{j}(z))$ at $z=\tau$ is $\nu$ as long as $r$ does not vanish. Since $z\mapsto\gamma z$ is a diffeomorphism, this coincides with the order of $p(z,\mathbf{j}(z))$ at $z=\gamma\tau$.
To conclude, since the map $\gamma\mapsto(\gamma\tau,c\tau+d,c)$ is injective on $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$, hence dominant on $\operatorname{\mathbb{C}}^{3}$, we only need to show that $r$ is non-trivial. Pick $\beta_{0}$ such that $\beta_{0}^{2\mu}=\alpha_{0}^{-1}$ and $U_{0}$ such that $U_{0}^{2\mu}=Y_{0}$. By Lemma˜6.1 at $\alpha^{2}=\mu\alpha_{0}$, $\beta=0$, $U_{1}^{2}=(\beta_{0}U_{0})^{-2(\mu-1)}Y_{1}$, there are $V_{1},V_{2}\in\operatorname{\mathbb{C}}(U_{1},U_{2})$ such that
@@
$p(\mathbf{j}(z))=\alpha^{2}(z-\gamma\rho)^{4}\left(6-\frac{2}{3}3^{2}+O(z-\gamma\rho)\right)=O((z-\gamma\rho)^{5}).$
@@
[…]
@@
[…]
Let $\beta$ be a root of $h(Y_{0})$ such that $\beta\notin\{0,1728\}$, and let $\tau$ be such that $j(\tau)=\beta$. In particular, $Y_{0}-\beta$ does not divide $p$. By Proposition˜6.2, $p(\gamma\tau,j(\gamma\tau))\neq 0$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ except for some proper Zariski closed set, and so $P$ has a pole at $\gamma\tau$, a contradiction. Therefore, $h(Y_{0})$ is of the form $\alpha Y_{0}^{s}(Y_{0}-1728)^{t}$ for some $\alpha\in\operatorname{\mathbb{C}}$.
Now $h(j(z))j^{\prime}(z)^{\ell}$ has order $3s+2\ell$ at all the conjugates of $\rho$ and $2t+\ell$ at all the conjugates of $i$, thus $p(j(z),j^{\prime}(z))$ must have at least the same order at those points. On writing $p=\sum_{\ell^{\prime}}Y_{1}^{\ell^{\prime}}p_{\ell^{\prime}}(Y_{0})$, Proposition˜6.6 implies that each $p_{\ell^{\prime}}(Y_{0})$ is divisible by $Y_{0}^{s+\lceil\frac{2(\ell-\ell^{\prime})}{3}\rceil}$ and by $(Y_{0}-1728)^{t+\lceil\frac{\ell-\ell^{\prime}}{2}\rceil}$. In particular, whenever $p_{\ell^{\prime}}\neq 0$ we have